OpenStudy (anonymous):

Radicals and stuff Please help last 2 questions !

OpenStudy (anonymous):

\[\sqrt[4]{9}\sqrt{9}/\sqrt[4]{9^5}\]

OpenStudy (jdoe0001):

\(\Large \bf \cfrac{\sqrt[4]{9}\sqrt{9}}{\sqrt[4]{9^5}}\quad ?\)

OpenStudy (anonymous):

Yes thats it

OpenStudy (anonymous):

@jdoe0001 can you help ? What you did there was the question I need help with solving

OpenStudy (jdoe0001):

ahemmm one sec

OpenStudy (anonymous):

Okay Thanks you

OpenStudy (jdoe0001):

\(\large {\bf \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}}\qquad \qquad a^{-\frac{{\color{blue} n}}{{\color{red} m}}} = \cfrac{1}{a^{\frac{{\color{blue} n}}{{\color{red} m}}}} \implies \cfrac{1}{\sqrt[{\color{red} m}]{a^{\color{blue} n}}} \\ \quad \\ \quad \\ \cfrac{\sqrt[4]{9}\sqrt{9}}{\sqrt[4]{9^5}}\implies\cfrac{\sqrt[4]{9}\sqrt{9^1}}{\sqrt[4]{9^5}}\implies \cfrac{9^{\frac{1}{4}}\cdot 9^{\frac{1}{2}}} {9^{\frac{1}{5}}}\implies \cfrac{9^{\frac{1}{4}}\cdot 9^{\frac{1}{2}}}{1}\cdot \cfrac{1}{9^{\frac{1}{5}}} \\ \quad \\ \cfrac{9^{\frac{1}{4}}\cdot 9^{\frac{1}{2}}}{1}\cdot 9^{-\frac{1}{5}} }\) so.. what do you think?

OpenStudy (jdoe0001):

recall your exponent rules -> http://www.math-play.com/image-exponents-rules.jpg <-- notice the 1st rule listed there

OpenStudy (jdoe0001):

hmmm shoot.. I just notice... the denominator has an exponent of 5....lemme fix that quick ahem so \(\large {\bf \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}}\qquad \qquad a^{-\frac{{\color{blue} n}}{{\color{red} m}}} = \cfrac{1}{a^{\frac{{\color{blue} n}}{{\color{red} m}}}} \implies \cfrac{1}{\sqrt[{\color{red} m}]{a^{\color{blue} n}}} \\ \quad \\ \quad \\ \cfrac{\sqrt[4]{9}\sqrt{9}}{\sqrt[4]{9^5}}\implies\cfrac{\sqrt[4]{9}\sqrt{9^1}}{\sqrt[4]{9^5}}\implies \cfrac{9^{\frac{1}{4}}\cdot 9^{\frac{1}{2}}} {9^{\frac{5}{4}}}\implies \cfrac{9^{\frac{1}{4}}\cdot 9^{\frac{1}{2}}}{1}\cdot \cfrac{1}{9^{\frac{5}{4}}} \\ \quad \\ \cfrac{9^{\frac{1}{4}}\cdot 9^{\frac{1}{2}}}{1}\cdot 9^{-\frac{5}{4}} }\)